dear sir, plz help me with the solution to the following problem A HOLLOW SPHERE OF RADIUS R IS FILLED BY WATER.THE TEMPERATURE OF ATMOSPHERE IS T WHICH IS BELOW 0 DEGREES.FIRSTLY ALL WATER COMES DOWN TO 0 DEGREES CENTIGRADE AND THEN STARTS FREEZING FROM THE OUTER SURFACE TOWARDS THE CENTRE.ASSUMING THAT THE THERMAL CONDUCTIVITY OF THE SHELL IS VERY HIGH, FIND THE TIME FOR THE WHOLE MASS OF WATER TO FREEZE.THERMAL CONDUCTIVITY OF ICE IS K AND LATENT HEAT OF FUSION IS L.

dear sir,


plz help me with the solution to the following problem


A HOLLOW SPHERE OF RADIUS R IS FILLED BY WATER.THE TEMPERATURE OF ATMOSPHERE IS T WHICH IS BELOW 0 DEGREES.FIRSTLY ALL WATER COMES DOWN TO 0 DEGREES CENTIGRADE AND THEN STARTS FREEZING FROM THE OUTER SURFACE TOWARDS THE CENTRE.ASSUMING THAT THE THERMAL CONDUCTIVITY OF THE SHELL IS VERY HIGH, FIND THE TIME FOR THE WHOLE MASS OF WATER TO FREEZE.THERMAL CONDUCTIVITY OF ICE IS K AND LATENT HEAT OF FUSION IS L.

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

dQ/dt=K(4(pi)r^2)T/(R-r)=Ldm/dt=Lρ(4p1(r^2)) dr/dt, Lρ(dr/dt)=KT/(R-r), ∫Lρ(R-r)dr=KT ∫dt